OpenStudy (vishweshshrimali5):

If a,b,c are real numbers such that: \(a^2 + b^2 + c^2 = 1\) prove the inequality: \(\cfrac{-1}{2} \le ab+bc+ca \le 1\)

OpenStudy (vishweshshrimali5):

@Miracrown @ganeshie8

OpenStudy (vishweshshrimali5):

Okay thanks @Miracrown I think I got that method also. Thanks for your help friends ..

Miracrown (miracrown):

You're welcome.

mathslover (mathslover):

@vishweshshrimali5 - I got the first part of the inequality : \((a+b+c)^2 \ge 0 \\ 1 + 2(ab + bc + ca) \ge 0 \quad ; \boxed{(a+b+c)^2 = a^2 + b^2 + c^2 + 2(ab + bc + ca) } \\ 2(ab + bc + ca) \ge - 1 \quad ;\space \\ ab + bc + ca \ge \cfrac{-1}{2} \\ \)

mathslover (mathslover):

Did you use AM-GM inequality to solve for the second inequality?

OpenStudy (vishweshshrimali5):

No

mathslover (mathslover):

Hmm, then ?

OpenStudy (vishweshshrimali5):

We used this: \((a-b)^2 + (b-c)^2 + (c-a)^2 \ge 0\)

mathslover (mathslover):

\(a^2 + b^2 - 2ab + b^2 + c^2 -2bc + c^2 + a^2 -2ca \ge 0\) \(2(a^2 + b^2 + c^2) -2(ab +bc+ca) \ge 0\) \(ab + bc + ca \le 1\) Cool! Got it. Thanks.

OpenStudy (vishweshshrimali5):

No problem.

mathslover (mathslover):

Though, I have another method for this.

mathslover (mathslover):

AM GM Inequality will also work here. Though, what you have used here, is also correct, but just wanted to put this up here : \[\cfrac{a^2 + b^2}{2} \ge \sqrt{a^2b^2} \\ \boxed{\color{Blue}{\cfrac{a^2 + b^2}{2} \ge ab}} ---------(1) \] \[\cfrac{b^2 + c^2}{2} \ge \sqrt{b^2 c^2} \\ \boxed{\color{blue}{\cfrac{b^2 + c^2}{2} \ge bc }} ---------(2) \] \[\cfrac{c^2 + a^2}{2} \ge \sqrt{c^2 a^2} \\ \boxed{\color{blue}{\cfrac{c^2 + a^2}{2} \ge ca }} ---------(3) \] Adding equations 1 , 2 and 3 : \[\cfrac{2(a^2 + b^2 + c^2)}{2} \ge ab + bc + ca \\ \cfrac{\cancel{2} (1)}{\cancel{2}} \ge ab + bc + ca \quad ; \boxed{a^2 + b^2 + c^2 = 1 } \space \text{(Given)} \\ 1 \ge ab + bc + ca \\ \text{OR} \quad \boxed{\color{red}{\bf{-\cfrac{1}{2} \ge ab + bc + ca\le 1}}} \]

OpenStudy (vishweshshrimali5):

Great work @mathslover

mathslover (mathslover):

Thank you! :-)